# 2Leibniz-Institut für Astrophysik Potsdam (AIP), An der Sternwarte 16, 14482 Potsdam, Germany

### v1

@@ Abstract
SDSS-V is carrying out a dedicated survey for white dwarfs, single and in binaries, and we report the analysis of the spectroscopy of 505 cataclysmic variables (CVs) and CV candidates obtained during the first 34 months of observations of SDSS-V. We developed a convolutional neural network (CNN) to aid with the identification of CV candidates among the over 2 million SDSS-V spectra obtained with the BOSS spectrograph. The CNN reduced the number of spectra that required visual inspection to $\simeq 2$ per cent of the total. We identified 779  CV spectra among the CNN-selected candidates, plus an additional 37  CV spectra that the CNN misclassified, but that were found serendipitously by human inspection of the data. Analysing the SDSS-Vspectroscopy and ancillary data of the 505  CVs in our sample, we report 62  new CVs, spectroscopically confirm 243 and refute 13  published CV candidates, and we report 68  new or improved orbital periods. We discuss the completeness and possible selection biases of the machine learning methodology, as well as the effectiveness of targeting CV candidates within SDSS-V. Finally, we re-assess the space density of CVs, and find $1.3\times 10^{-5}\,\mathrm{pc^{-3}}$.
@@ 1 Introduction
The Sloan Digital Sky Survey (SDSS – York et al. 2000; Gunn et al. 2006) has revolutionised the process of identifying and classifying CVs by providing large-scale spectroscopic data. These spectra have yielded many short-period and non-outbursting CVs (Szkody et al., 2002, 2003a, 2004, 2005, 2006, 2007, 2009, 2011; Inight et al., 2023a, b) and helped to address the selection biases within the historic CV samples (Gänsicke et al., 2009). All of these spectra were manually identified using visual inspection. Following 20 years of using hard-wired optical fibre configurations called “plug plates” (Blanton et al., 2003), SDSS-V  (Kollmeier et al., 2017) now uses a robotic fibre positioning system (Jurgenson et al., 2020) and the fibre-fed optical BOSS spectrograph (Smee et al., 2013) to capture up to 500 spectra simultaneously resulting in the 2 008 677 spectra [^1] In more detail, the data set we used is the full set of SDSS-V spectra reduced with the BOSS pipeline version v6_1_0, excluding 1 191 969 spectra where the robot was placed to feed the infrared APOGEE spectrograph. Those fibres record in most cases a blank sky spectrum, or, rarely, that of a random object falling into the fibre. These “serendipituous” BOSS spectra are easily identified by their file name structure, spec-{FIELD}-{MJD}-{RA}{DEC}.fits obtained since the beginning of SDSS-V in December 2020 up to 2023 September 9.
SDSS-V was also the first spectroscopic survey to specifically include CV candidates as targets (Inight et al., 2023b) – accounting for 117 841  spectra out of the total 2 008 677. This has proved invaluable, both for confirming existing CV candidates and also uncovering new CVs identified by an ultraviolet excess. However we recall that CVs discovered by SDSS in the first 20 years were serendipitously identified while searching for other types of objects, such as quasars. A comprehensive identification of CVs therefore requires an analysis of the entire set of spectra in SDSS-V. A manual inspection of all spectra would be a time-consuming and error-prone exercise. We therefore take an automated approach to reduce the number of spectra requiring expert review. This approach will also be useful for analysing other, even more prolific, spectroscopic surveys such as DESI (DESI Collaboration et al., 2016a, b), WEAVE (Dalton et al., 2012) and 4MOST (de Jong et al., 2016).
@@ 2 Machine learning to classify large survey spectroscopy
We describe here our approach for shortlisting the spectra of potential CVs. Once a shortlist has been generated, we visually inspect the spectra and obtain orbital periods using the methods described in Inight et al. (2023b). We therefore do not discuss classification and orbital periods any further here and focus upon the initial generation of shortlists.
Machine learning has been used to find CVs from survey data before. Mistry et al. (2022, 2024) used machine learning to analyse photometry from Gaia Alerts and ZTF, respectively, yielding 2 800similar-to-or-equalsabsent2800$\simeq$ 2\,800≃ 2 800 CVs from 13 2801328013\,28013 280 Gaia Alerts. Hou et al. (2020) analysed specific emission and absorption lines in several million LAMOST spectra and found 380 CV spectra. Vincent et al. (2023) used machine learning based upon neural network technology to identify and classify white dwarfs. More generally Echeverry et al. (2022); Vincent et al. (2024) and García-Zamora et al. (2023) used a binary machine learning classifier to analyse Gaia spectra in the search for white dwarfs.
@@
Table 1: Summary of the spectroscopic samples used to train and test the CNN. The initial samples were constructed as detailed in Section, 2.1.12.1.5 and then randomly sub-sampled. see Section 3.1 for details.
Initial Final
CVs 623 563*
Galaxies 50 000 1206
@@
Detached white dwarf plus main sequence binary 1602 1198
@@ 2.1.3 Quasars
We took our quasar sample from Schneider et al. (2003). Although there are later, larger catalogues we considered that that this early work would be particularly reliable, having been subject to a number of manual examinations. The table was downloaded from TapVizieR (Landais et al., 2013) using the following statement:
@@ 2.1.4 Stars
. The SDSS pipeline attributes a spectral type (based on the subClass field in the SpecPhotoAll table) to each of these stars. There were 34 different subClass values present and a random sample of 35 of each was selected. In some cases there were fewer than 35 for a given subClass and hence the stellar sample consists of only 842 spectra.
@@ 2.1.6 Detached white dwarf plus main sequence binaries
Figure 1: Frequency distribution of the individual CVs in the train/test sample (Table 1) being identified as a false negative by the CNN, from 200 random permutations of the sample. The horizontal axis represents the individual 623 CVs ordered by frequency. 220 CVs were always identified correctly and 90 per cent of the CVs occur at most three times as a false negative. 60 CVs frequently resulted as false negatives indicating that they are not suitable for use in the training data (Section 2.1.1.)
@@ 3.1 Training
The CNN is trained by splitting the train/test dataset into two randomly chosen samples with 80 per cent being used for training and the remaining 20 per cent being tested to evaluate the result. To explore the effect of the random selection of the test sample we repeated the training 200 times with the selection of the training and test samples being determined randomly. Whilst the overall numbers of false positives and false negatives did not vary significantly it became apparent that certain CVs were consistently presenting as false negatives. We therefore analysed the set of false negatives in the 200 CNNs (see Fig. 1). From this it became apparent that 60 of the CVs were rarely if ever being found when included in the test sample. A review of the spectra of these 60 CVs revealed that 13 had significant quality issues (due to low S/N or a catastrophic failure in the pipeline), 11 exhibited cyclotron humps, 11 had relatively featureless blue novalike spectra, and nine were obtained during dwarf nova outbursts and resemble hot stars 333 The remaining 16 were not exceptional.. It is therefore not surprising that these 60 CV spectra were not correctly identified by the CNN. As these 60 CVs were not usefully contributing to the training data we removed them from the train/test dataset. The final CNN was then computed based on this revised train/test dataset of 563 CVs. The potential bias introduced by this is discussed further in Section 4.3.3.
The final CNN was trained on the train/test dataset listed in Table 1. We computed a Confusion Matrix for the final model (Fig. 2) which shows that only six false positives were predicted and $110$ (out of a potential $113$) true positives were found.
@@
CV-related cartons (4.1) 117 841 688 1054
All SDSSV (4.2) 2 008 677 779 38402
@@ 3.2.1 The plate survey of SDSS-V
A reliable set of 118 CVs were identified from the first eight months of SDSS-V (the “plate” survey, Inight et al. 2023b). There are 302  observations of these 118 CVs amongst the 33 064 BOSS spectra targeted by the CV-related cartons444mwm_wd_core, mwm_cb_gaiagalex, mwm_cb_uvex1, mwm_cb_uvex2, mwm_cb_uvex3, mwm_cb_uvex4 and mwm_cb_uvex5, mwm_cb_cvcandidates (see Almeida et al. 2023 section 9.2). We analysed these 33 064 spectra which were reduced with the SDSS-V pipeline version 6_1_0 (Almeida et al., 2023) using our CNN resulting in 254  out of 302 CV spectra being recovered with only 439  false positives (Table  2). Whilst 118 spectra were used in the train/test dataset the remaining 184  were unseen data. The low level ($\simeq 1$ per cent) of false positives is reassuring.
@@ 3.2.2 Previously known CV candidates
The mwm_cb_cvcandidates carton of SDSS-V was generated from lists of candidate or known CVs and was therefore expected to contain a higher proportion of CVs than other cartons. Our CNN identified 200 of the 288 spectra of targets in mwm_cb_cvcandidates obtained by SDSS-V  as CVs. A subsequent human review of the other 88 spectra found that 35 were CVs but with poor spectrum quality, 38 were not CVs (i.e the CNN correctly identified them as true negatives), and only 15 were spectra of actual CVs with good quality spectra. The CVs identified from this are included in the supplementary data.
@@ 3.2.3 The BOSS QA sample
We reviewed the sample of candidate CVs provided by the QA team, removed non-CVs and used the remainder as an independent check on the completeness of our CNN. Our CNN identified 329  out of the 366  confirmed CVs identified by the QA team.
@@ 3.2.4 All spectra in SDSS I to IV
The vast majority of CVs identified among the spectra obtained within SDSS I to IV were found via visual inspection (Szkody et al. 2002, 2003a, 2004, 2005, 2006, 2007, 2009, 2011, see Inight et al. 2023a for a detailed overview). A grand total of 507  CVs (687  spectra) for which SDSS spectroscopy is available, with classifications at the sub-type level for most of them (Inight et al., 2023a), provides the largest sample to test the CNN. We analysed all 5 764 653 BOSS spectra obtained by SDSS I to IV with our CNN, which recovered 616  out of 687  spectra representing 471  of the 507  known CVs (from Inight et al. 2023a). Whilst 507  spectra were used for training, 180   spectra were unseen by the CNN.
@@ 4.1 CV-related cartons in SDSS-V
Up to $\mathrm{MJD}=60196$, SDSS-V obtained 117 841  BOSS spectra of targets within one or more CV-related cartons. This sample included 295 spectra of previously known CVs plus an unknown number of new CVs. As expected our CNN recovered $254$ out of the $295$ CV spectra and found an additional 1488 candidate spectra. An analysis of the 19 CVs (41 spectra)
that were not recovered showed that 14 had severe flux calibration issues, whilst the remainder had indistinct emission lines. The 1488 spectra were manually examined and 434 were found to be CV spectra yielding a total of 688 true positives. The CVs identified from this examination are included in the supplementary data.
@@ 4.2 All targets in SDSS-V
Recalling that some CVs in SDSS-V will only be found in unrelated cartons, we analysed the entire 2 008 677 spectra in SDSS-V up to $\mathrm{MJD}=60196$ using our CNN resulting in 39 181 spectra ($\simeq 2$ per cent of the total) being classified as candidate CVs. These spectra were examined by eye for CV-related features yielding 779 true positives and 38 402 false positives. Any CVs identified from this are included in the supplementary data.
@@ 4.3.1 False positives
To characterise the false positives arising from the CNN we analysed the 1488 spectra identified in Section 4.1. 1054 of these spectra proved not to be CVs. Of these 1054 spectra 99 were catastrophic failures with random profiles. All but three of the remaining 955 spectra exhibited some kind of emission lines. These were all thin lines unlike those of CVs. 474 of these 955 included H  or He  emission lines.
@@ 4.3.2 Does the S/N of spectra affect the reliability of the CNN?
Before discussing selection biases of individual CV sub-types we consider whether low S/N systematically affects the outcome of our CNN analysis [^5] It is important to note that there are reasons other than low S/N for a poor quality spectrum, including poor flux calibration, partially missing spectral coverage, or strong sky line residuals. We analysed the effect of S/N by considering the rate of false positives in the sample of 2 008 677  spectra in SDSS-V up to $\mathrm{MJD}=60196$ (Fig. 3). To assess the false negatives was impractical as it would require the manual analysis of all 2 008 677 spectra and we assume that the behaviour of the false negative rate is similar to that of the false positive rate.
Figure 3: The fraction of false positives with respect to the sum of false plus true positives as a function of S/N based on the 39 181 CV candidates that the CNN identified among the total 2 008 677 spectra that were analysed. The proportion of false positives is relatively flat, $\simeq 0.95$, until $\mathrm{S/N}\lesssim 1$. Nevertheless, even at very low S/N some true positives are identified. The relatively small number of spectra with $S/N\gtrapprox 15$ yield relatively few false positives.

### v2

@@ Abstract
SDSS-V is carrying out a dedicated survey for white dwarfs, single and in binaries, and we report the analysis of the spectroscopy of 504 cataclysmic variables (CVs) and CV candidates obtained during the first 34 months of observations of SDSS-V. We developed a convolutional neural network (CNN) to aid with the identification of CV candidates among the over 2 million SDSS-V spectra obtained with the BOSS spectrograph. The CNN reduced the number of spectra that required visual inspection to $\simeq 2$ per cent of the total. We identified 776  CV spectra among the CNN-selected candidates, plus an additional 27  CV spectra that the CNN misclassified, but that were found serendipitously by human inspection of the data. Analysing the SDSS-V spectroscopy and ancillary data of the 504  CVs in our sample, we report 61  new CVs, spectroscopically confirm 248 and refute 13  published CV candidates, and we report 82  new or improved orbital periods. We discuss the completeness and possible selection biases of the machine learning methodology, as well as the effectiveness of targeting CV candidates within SDSS-V. Finally, we re-assess the space density of CVs, and find $1.2\times 10^{-5}\,\mathrm{pc^{-3}}$.
@@ 1 Introduction
The Sloan Digital Sky Survey (SDSS – York et al. 2000; Gunn et al. 2006) has revolutionised the process of identifying and classifying CVs by providing large-scale spectroscopic data. These spectra have yielded many short-period and non-outbursting CVs (Szkody et al., 2002, 2003a, 2004, 2005, 2006, 2007, 2009, 2011; Inight et al., 2023a, b) and helped to address the selection biases within the historic CV samples (Gänsicke et al., 2009). All of these spectra were manually identified using visual inspection. Following 20 years of using hard-wired optical fibre configurations called “plug plates” (Blanton et al., 2003), SDSS-V  (Kollmeier et al., 2017) now uses a robotic fibre positioning system (Jurgenson et al., 2020) and the fibre-fed optical BOSS spectrograph (Smee et al., 2013) to capture up to 500 spectra simultaneously resulting in the 2 008 677 spectra [^1] In more detail, the data set we used is the full set of SDSS-V spectra reduced with the BOSS pipeline version v6_1_0, excluding 1 191 969 spectra where the robot was placed to feed the infrared APOGEE spectrograph. Those fibres record in most cases a blank sky spectrum, or, rarely, that of a random object falling into the fibre. These “serendipituous” BOSS spectra are easily identified by their file name structure, spec-{FIELD}-{MJD}-{RA}{DEC}.fits obtained since the beginning of SDSS-V on 2020 October 24 up to 2023 September 9.
SDSS-V was also the first spectroscopic survey to specifically include CV candidates as targets (Inight et al., 2023b) – accounting for 117 841  spectra out of the total 2 008 677. This has proved invaluable, both for confirming existing CV candidates and also uncovering new CVs targeted because of an ultraviolet excess. However we recall that CVs discovered by SDSS in the first 20 years were serendipitously identified while searching for other types of objects, primarily quasars. A comprehensive identification of CVs therefore requires an analysis of the entire set of spectra in SDSS-V. A visual inspection of all spectra would be a time-consuming and error-prone exercise. We therefore developed an automated approach to reduce the number of spectra requiring expert review. This approach will also be useful for analysing other, even more prolific, spectroscopic surveys such as DESI (DESI Collaboration et al., 2016a, b), WEAVE (Dalton et al., 2012) and 4MOST (de Jong et al., 2016).
@@ 2 Machine learning to classify large survey spectroscopy
We describe here our approach for shortlisting the spectra of potential CVs. Once a shortlist has been generated, we visually inspect the spectra and obtain orbital periods using the methods described in Inight et al. (2023b). We therefore do not discuss details on determining classifications and orbital periods any further here and focus upon the initial generation of shortlists.
Machine learning has been used to find CVs from survey data before. Mistry et al. (2022, 2024) used machine learning to analyse photometry from Gaia Alerts and ZTF, respectively, yielding 2 800similar-to-or-equalsabsent2800$\simeq$ 2\,800≃ 2 800 CVs candidates from 13 2801328013\,28013 280 Gaia Alerts. Hou et al. (2020) analysed specific emission and absorption lines in several million LAMOST spectra and found 380 CV spectra. Vincent et al. (2023) used machine learning based upon neural network technology to identify and classify white dwarfs. More generally Echeverry et al. (2022); Vincent et al. (2024) and García-Zamora et al. (2023) used a binary machine learning classifier to analyse Gaia spectra in the search for white dwarfs.
@@
Table 1: Summary of the spectroscopic samples used to train and test the CNN. The initial samples were constructed as detailed in Sections, 2.1.12.1.5 and then randomly sub-sampled to create the final test/train dataset. The test data is a randomly selected subset of 20 per cent of the final test/train dataset and the size of the test data is shown here for ease of comparison with Fig. 2.
see Section 3.1 for details.
Initial Final Test
CVs 623 563* 113
[…]
@@
Detached white dwarf plus 1602 1198
            main sequence binary
@@ 2.1.3 Quasars
We took our quasar sample from Schneider et al. (2003). Although there are later, larger catalogues we considered that this early work would be particularly reliable, having been subject to a number of visual examinations. The table was downloaded from TapVizieR (Landais et al., 2013) using the following statement:
@@ 2.1.4 Stars
The SDSS pipeline attributes a spectral type (based on the subClass field in the SpecPhotoAll table) to each of these stars. There were 34 different subClass values present and a random sample of 35 of each was selected. In some cases there were fewer than 35 for a given subClass and hence the stellar sample consists of only 842 spectra.
@@ 2.1.6 Detached white dwarf plus main sequence binaries
Figure 1: Frequency distribution of the individual CVs in the train/test sample (Table 1) being identified as a false negative by the CNN, from 200 random permutations of the sample. The horizontal axis represents the individual 623 CVs ordered by frequency. 220 CVs were always identified correctly and 90 per cent of the CVs occur at most three times as a false negative. 60 CVs frequently ($f>6$) resulted as false negatives indicating that they are not suitable for use in the training data (Section 2.1.1.)
@@ 3.1 Training
The CNN is trained by splitting the train/test dataset into two randomly chosen samples with 80 per cent being used for training and the remaining 20 per cent being tested to evaluate the result. To explore the effect of the random selection of the test sample we repeated the training 200 times with the selection of the training and test samples being determined randomly. Whilst the overall numbers of false positives and false negatives did not vary significantly it became apparent that certain CVs were consistently presenting as false negatives. We therefore analysed the set of false negatives in the 200 CNNs (see Fig. 1). From this it became apparent that 60 of the CVs were rarely if ever being found when included in the test sample. A review of the spectra of these 60 CVs revealed that 13 had significant quality issues (due to low S/N or a catastrophic failure in the pipeline), 11 exhibited cyclotron humps, 11 had relatively featureless blue novalike spectra, and nine were obtained during dwarf nova outbursts and resemble hot stars 333 The remaining 16 were not exceptional.. It is therefore not surprising that these 60 CV spectra (see Table 3 in the supplementary data) were not correctly identified by the CNN. As these 60 CVs were not usefully contributing to the training data we removed them from the train/test dataset. The final CNN was then computed based on this revised train/test dataset of 563 CVs. The potential bias introduced by this is discussed further in Section 4.3.3.
The final CNN was trained using the train/test dataset listed in Table 1. The training used 80 per cent of the CVs from the train/test database and 80 per cent of the non-CVs. The remaining 20 per cent of CVs and non-CVs were used to test the model (see final column in Table 1. We then computed a Confusion Matrix for the final model (Fig. 2) which shows that only six false positives were predicted and $110$ (out of a potential $113$) true positives were found.
@@
CV-related cartons (4.1) 117 841 685 1054
All SDSSV (4.2) 2 008 677 776 38 402
@@ 3.2.1 The plate survey of SDSS-V
A reliable set of 118 CVs were identified from the first eight months of SDSS-V (the “plate” survey, Inight et al. 2023b). There are 302  observations of these 118 CVs amongst the 33 064 BOSS spectra targeted by the CV-related cartons444mwm_wd_core, mwm_cb_gaiagalex, mwm_cb_uvex1, mwm_cb_uvex2, mwm_cb_uvex3, mwm_cb_uvex4 and mwm_cb_uvex5, mwm_cb_cvcandidates (see Almeida et al. 2023 section 9.2). We analysed these 33 064 spectra which were reduced with the SDSS-V pipeline version 6_1_0 (Almeida et al., 2023) using our CNN resulting in 254  out of 302 CV spectra (99 CVs out of 118 - 84 per cent) being recovered with only 439  false positives (Table  2). Whilst 118 spectra were used in the train/test dataset the remaining 184  were unseen data. The low level ($\simeq 1$ per cent) of false positives is reassuring.
@@ 3.2.2 Previously known CV candidates
The mwm_cb_cvcandidates carton of SDSS-V was generated from lists of candidate or known CVs and was therefore expected to contain a higher proportion of CVs than other cartons. Our CNN identified 200 of the 288 spectra of targets in mwm_cb_cvcandidates obtained by SDSS-V  as CVs. A subsequent human review of the other 88 spectra found that 35 were CVs but with poor spectrum quality, 38 were not CVs (i.e the CNN correctly identified them as true negatives), and only 15 were spectra of actual CVs with good quality spectra. The CVs identified from this are included in Table 1 in the supplementary data.
@@ 3.2.3 The BOSS QA sample
We reviewed the sample of candidate CV spectra provided by the QA team, removed non-CVs and used the remainder as an independent check on the completeness of our CNN. Our CNN identified 329  out of the 366  (90 per cent) confirmed CV spectra identified by the QA team. It is important to note that in some cases the CNN will have detected the CV from a different spectrum and hence the recovery rate will be slightly understated.
@@ 3.2.4 All spectra in SDSS I to IV
The vast majority of CVs identified among the spectra obtained within SDSS I to IV were found via visual inspection (Szkody et al. 2002, 2003a, 2004, 2005, 2006, 2007, 2009, 2011, see Inight et al. 2023a for a detailed overview). A grand total of 507  CVs (687  spectra) for which SDSS spectroscopy is available, with classifications at the sub-type level for most of them (Inight et al., 2023a), provides the largest sample to test the CNN. We analysed all 5 764 653 BOSS spectra obtained by SDSS I to IV with our CNN, which recovered 616  out of 687  spectra representing 471  (93 per cent) of the 507   known CVs (from Inight et al. 2023a). Whilst 507  spectra were used for training, 180   spectra were unseen by the CNN.
@@ 4.1 CV-related cartons in SDSS-V
Up to $\mathrm{MJD}=60196$, SDSS-V obtained 117 841  BOSS spectra of targets within one or more CV-related cartons. This sample included 295 spectra of previously known CVs plus an unknown number of new CVs. As expected our CNN recovered $254$ out of the $295$ CV spectra and found an additional 1488 candidate spectra. An analysis of the 19 CVs (41 spectra) that were not recovered showed that 14 had severe flux calibration issues, whilst the remainder had indistinct emission lines. The 1488 spectra were manually examined and 431 were found to be CV spectra yielding a total of 685   true positives (422  CVs).
@@ 4.2 All targets in SDSS-V
Recalling that some CVs in SDSS-V will only be found in unrelated cartons, we analysed the entire 2 008 677 spectra in SDSS-V up to $\mathrm{MJD}=60196$ using our CNN resulting in 39 181 spectra ($\simeq 2$ per cent of the total) being classified as candidate CVs. These spectra were examined by eye  [^5] The co-authors reviewed randomly selected subsets of the 39 181 and then the author separately reviewed the entire set to ensure consistency. The process was facilitated by some bespoke python code so that each spectrum could be classified using a single keystroke. for CV-related features yielding 776 true positives and 38 402 false positives. The 776 spectra include 55 CVs that were not in the CV-related cartons.
@@ 4.3.1 False positives
To characterise the false positives arising from the CNN we analysed the 1488 spectra identified in Section 4.1. 1054 of these spectra proved not to be CVs. Of these 1054 spectra 99 were catastrophic failures with random appearance. All but three of the remaining 955 spectra exhibited some kind of emission lines. These were all thin lines unlike those of CVs. 474 of these 955 included H  or He  emission lines.
@@ 4.3.2 Does the S/N of spectra affect the reliability of the CNN?
Before discussing selection biases of individual CV sub-types we consider whether low S/N systematically affects the outcome of our CNN analysis [^6] It is important to note that there are reasons other than low S/N for a poor quality spectrum, including poor flux calibration, partially missing spectral coverage, or strong sky line residuals. We analysed the effect of S/N by considering the fraction of false positives in the sample of 2 008 677  spectra in SDSS-V up to $\mathrm{MJD}=60196$ (Fig. 3). To assess the false negatives was impractical as it would require the visual analysis of all 2 008 677 spectra and we assume that the behaviour of the false negative rate is similar to that of the false positive rate. The proportion of false positives ($\simeq 0.95$) is relatively flat for $1<\mathrm{S/N}<10)$.
Figure 3: The fraction of false positives with respect to the sum of false plus true positives as a function of S/N based on the 39 181 CV candidates that the CNN identified among the total 2 008 677 spectra that were analysed. The proportion of false positives is relatively flat, $\simeq 0.95$, until $\mathrm{S/N}\lesssim 1$. Nevertheless, even at very low S/N some true positives are identified. The relatively small number of spectra with $\mathrm{S/N}\gtrapprox 15$ yield relatively few false positives.